A uniform electric field of magnitude 6.6 N/C passes through a circle of radius 13 cm.

A) What is the electric flux through the circle when its face is perpendicular to the field lines? Express your answer using two significant figures.

B) What is the electric flux through the circle when its face is at 45∘ to the field lines? Express your answer using two significant figures.

C) What is the electric flux through the circle when its face is parallel to the field lines? Express your answer using two significant figures.
Answer #1

The magnitude of the dot product E⋅A = E*A*cosθ where θ is the angle between the flux lines and a perpendicular to the area.
A ) when the electric feild is perpendiculer to the circle surface than the angle between these two will 0°.

So flux =
E*A*cosθ = 6.6*π*0.13²*cos0 = 0.3504 N⋅m²/C <------- A
B.) When angle is 45° than electric flux
= 6.6*π*0.13²*cos45°= 0.2478 N⋅m²/C <------- B
C.) When electric field faces perpendicularlly than it will make 90° angle. so flux
= 6.6*π*0.13²*cos90 = 0 N⋅m²/C <------- C